OpenStudy (anonymous):

VERTICAL ASYMPTOTES AND CHECK MY WORK y=cot(x-pi)+1 a. list two vertical asymptotes-(help me) b. Period-pi c. phase shift- pi to the right d. vertical translation- up 1

OpenStudy (anonymous):

is all answers are right

OpenStudy (anonymous):

yes all are true

OpenStudy (anonymous):

I need help with finding the vertical asymptotes pretty please

OpenStudy (tkhunny):

What are the zeros of \(tan(x-\pi)\)? \(\cot(x-\pi) = \cos(x-\pi)/\sin(x-\pi)\). Where is \(\sin(x-\pi) = 0\)?

OpenStudy (anonymous):

at 0 degrees and 180 degrees and 360 degrees

OpenStudy (tkhunny):

\(\sin(x-\pi) = \sin(x)\cos(\pi) - \cos(x)\sin(\pi) = -\sin(x) - 0 = -\sin(x)\) Where is \(\sin(x-\pi) = 0\) Same Question: Where is \(-\sin(x) = 0\) Sure enough, \(x\in \{0,\pi,2\pi\}\) or 0º, 180º, 360º if you like. There are your vertical asymptotes.

OpenStudy (anonymous):

thank you thank you